The equation is h(t)= -16t^2+80t. (h for feet and t for seconds). The question asks how much time it takes for the ball to reach 96 feet. I tried plugging 96 in for h(t) and then subtracting it and inserting it on the other side. I factored, but got t=3 and t=2. I don't think I am supposed to get two positive numbers.
Found 2 solutions by richwmiller, Alan3354:
Answer by richwmiller(17219)

The ball goes up and then comes down so there will be two times at that height unless 1) it never reaches that height or 2) it reaches exactly that height and no more
Your answers are not correct.
16t^2-80t+92 = 0
4t^2-20t+23=0
which doesn't factor
1.8 and 3.2 are closer

Answer by Alan3354(69443)

The equation is h(t)= -16t^2+80t. (h for feet and t for seconds). The question asks how much time it takes for the ball to reach 96 feet. I tried plugging 96 in for h(t) and then subtracting it and inserting it on the other side. I factored, but got t=3 and t=2. I don't think I am supposed to get two positive numbers.
---------------
Check your answers.
h(t)= -16t^2+80t
h(2) = -16*4 + 80*2 = -64 + 160
h(2) = 96 That's correct.
------------
h(3) = -16*9 + 80*3 = -144 + 2400
h(3) = 96 Also correct.
------------
It goes up thru 96', then it comes down.